# supReferences

### v4

@@ 4.3 Discussion
Due to the presence of the $\log(N)$ term in the definition of $\delta_{N}$, the dependence in $N$ of bound given in Theorem 2 for $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ is slightly worse than that of the probabilistic bound obtained in Theorem 1 for PFs. On the other hand, in the context of this work, using RQMC within PF algorithms has the advantage to produce a non-linear filleting algorithm having time uniform and almost sure guarantees. In particular, it follows from Theorem 2 that if $N\in\{2^{k},\,k\in\mathbb{N}\}$ and, for some $\kappa\in(0,1)$, we have
$N\geq\frac{\log(N)+3}{2}\inf_{\gamma\in(0,1)}\kappa^{-\frac{2}{\gamma}}\bigg{(}\frac{\bar{C}_{2}}{1-\epsilon_{\star}^{1-\gamma}}\bigg{)}^{\frac{2}{\gamma}}$ (10)
then $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ with $\mathbb{P}$-probability one.
@@
Unfortunately, we cannot compare the lower bound on $N$ given in (10) with the one obtained in (7), used to ensure that with probability at least $1-q$ the PF estimates $\{\hat{\eta}_{t}^{N}\}_{t=1}^{T}$ are such that $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$, the reason being that there is no reason for the constants $\bar{C}_{1}$ and $\bar{C}_{2}$ appearing in these bounds to be identical. Given the properties of scrambled $(t,s)$-sequences, reminded in Section 4.1, it is however reasonable to expect that, in the worst case, the constant $\bar{C}_{2}$ in (10) can only be slightly larger than the constant $\bar{C}_{1}$ appearing in (7). In any cases, if we want to compute estimates $\{\hat{\eta}_{t}^{N}\}_{t=1}^{T}$ of $\{\hat{\eta}_{t}\}_{t=1}^{T}$ such that $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ with probability at least $1-q$, then it follows from Theorem 1 and Theorem 2 that there exists a value $T_{\kappa,q}\in\mathbb{N}$ such that SQMC is computationally cheaper than PFs for performing this task for any $T\geq T_{\kappa,q}$. Based on (7) and (10), it is reasonable to expect that $T_{\kappa,q}$ is an increasing function of both $\kappa$ and $q$.
The conclusions drawn in this subsection are based on a toy filtering problem, and future research should aim at establishing if they remain valid when considering more realistic filtering tasks.
5 Concluding remarks
Given the popularity of PF algorithms in many applications it is important to understand the behaviour of the resulting joint estimation error, $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$. Surprisingly, with the exception of Del Moral, (2013, Corollary 14.5.7, page 449), this latter quantity (or a variant of it) does not seem to have been previously studied in the literature. The results we obtained in this paper are reassuring, but an important question that has yet to be studied is how the dimension of the filtering problem impacts the behaviour of $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.
Lemma 1, the key result that allows us to study $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ for PFs as well as the quantity $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ for SQMC, is obtained for a toy filtering problem but we expect that it can be generalized by following a similar proof technique. The proof of this lemma, which studies the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ for a given realization $(u_{i})_{i\geq 1}$ of the $\mathcal{U}(0,1)$ random numbers underpinning the algorithm, relies notably on three key building blocks, namely a result showing that the SSM forgets exponentially quickly its initial distribution $\eta_{1}$ as $t\rightarrow\infty$ (illustrated by Proposition 1), a bound for $\big{\|}\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}-\hat{\eta}_{t-1}^{N}M\big{\|}$ (Lemma 8 in the appendix) and a stability result for the particles generated by the algorithm (Lemma 9 in the appendix). The first of these three building blocks is standard, in the sense that in the literature on PFs all the existing bounds for $\sup_{t\geq 1}\mathbb{E}[|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|]$ have been established under the assumption that the model forgets exponentially quickly its initial distribution (see Caffarel et al.,, 2025, and references therein), an assumption that has been shown to hold for a large class of SSMs (see e.g. Del Moral et al.,, 2023). By contrast, to prove Lemma 1 we derive a bound for $\big{\|}\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}-\hat{\eta}_{t-1}^{N}M\big{\|}$ which applies only to some one-dimensional SSMs and a stability result for the particle system which is tailored to the specific filtering problem addressed in this work. We stress that this latter result is quite strong, as it shows (under a reasonable assumption on $(u_{i})_{i\geq 1}$) that there exists a bounded interval $I\subset\mathbb{R}$ that contains, at any time instant $t\geq 1$, at least a proportion $\varsigma>0$ of the particles. Our conjecture is that a result similar to that of Lemma 1 can be established for more realistic filtering problems where this strong stability property of the particle system cannot hold. For instance, we expect that this is the case for the SSM considered in this paper if, instead of assuming that $y_{t}=0$ for all $t\geq 1$, we assume that the sequence of observations $(y_{t})_{t\geq 1}$ is unbounded but such that $\sup_{t\geq 1}|y_{t+1}-y_{t}|<\infty$.
In this work, we focussed on the use of scrambled $(t,s)$-sequences within PF algorithms to improve the safety of the filtering procedure, that is to obtain filtering algorithms for which we have $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=0$ with $\mathbb{P}$-probability one. Future research should aim at determining if other approaches can be used to derive safe filtering algorithms.
Appendix A Proofs
A.1 Additional notation
@@
For a function $f\in\mathcal{C}^{1}(\mathbb{R})$ we let $V(f)=\int_{\mathbb{R}}|f^{\prime}(x)|\mathrm{d}x$ with the convention that $V(f)=0$ if $f^{\prime}(x)=0$ for all $x\in\mathbb{R}$, and, abusing notation, for a function $f:\mathbb{R}\rightarrow\mathbb{R}$ continuously differentiable on some non-empty interval $(a,b)$ we let $V(\mathbbm{1}_{(a,b]}f)=V(\mathbbm{1}_{(a,b)}f)=\int_{a}^{b}|f^{\prime}(x)|\mathrm{d}x$. For all $(z,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $M_{r,s^{2}}(z,\mathrm{d}x)=\mathcal{N}_{1}(rz,s^{2})$ and $G_{s}(z)=e^{-sz^{2}/2}$, and for any bounded function $H:\mathbb{R}\rightarrow(0,\infty)$ we let $\Psi_{H}:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be defined by
@@
Finally, we let $\varphi(\cdot)$ and $\Phi(\cdot)$ denote the probability and cumulative density function of the $\mathcal{N}_{1}(0,1)$ distribution, respectively, and for all $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $h_{a,r,s}:\mathbb{R}\rightarrow[-1,1]$ be defined by
$h_{a,r,s}(x)=\Phi\bigg{(}\frac{a-rx}{s}\bigg{)}-\Phi\bigg{(}\frac{a}{s}\bigg{)},\quad x\in\mathbb{R}.$ (11)
A.2 Preliminary results
A.2.1 A simple but useful technical lemma

### v5

@@ 3.3.1 Some remarks
Due to the presence of the $\log(N)$ term in the definition of $\delta_{N}$, the dependence in $N$ of the bound given in Theorem 2 for $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ is slightly worse than that of the probabilistic bound obtained in Theorem 1 for PFs. On the other hand, in the context of this work, SQMC has the advantage to be a filtering algorithm having time uniform and almost sure guarantees.
@@
Unfortunately, we cannot compare the lower bound on $N$ given in Corollary 2 with the one obtained in Corollary 1 to ensure that, with probability at least 1q1-q, the PF estimates $\{\hat{\eta}_{t}^{N}\}_{t=1}^{T}$ are such that $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$: there is indeed no reason for the constants $\bar{C}_{1}$ and $\bar{C}_{2}$ appearing in these bounds to be identical. Given the properties of scrambled $(t,s)$-sequences reminded in Section 1.4, it is however reasonable to expect that, in the worst scenario, the constant $\bar{C}_{2}$ appearing in Corollary 2 can only be slightly larger than the constant $\bar{C}_{1}$ appearing in Corollary 1.
4 Proof of Theorems 1 and 2 and of Corollaries 1 and 2
4.1 Preliminary result: A general discrepancy bound for filtering algorithms
Let $\Phi,\Psi:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be defined
$\Psi(\mu)(\mathrm{d}x)=\frac{G(x)\mu(\mathrm{d}x)}{\mu(G)},\quad\Phi(\mu)=\Psi(\mu)M,\quad\mu\in\mathcal{P}(\mathbb{R})$
and note that $\eta_{t+1}=\Phi(\eta_{t})$ for all $t\geq 1$.
Using this notation, the following result plays a central role in the proof of Theorems 1 and 2:
Lemma 1.
There exists a constant $C\in(0,\infty)$ such that, for any sequence $(\mu_{t})_{t\geq 1}$ in $\mathcal{P}(\mathbb{R})$ and letting $\hat{\mu}_{t}=\Psi(\mu_{t})$ for all $t\geq 1$, we have, for any $T\geq 1$,
$\sup_{t\in\{1,\dots,T\}}\|\hat{\mu}_{t}-\hat{\eta}_{t}\|\leq C\,\Delta_{T}\log\big(1+\Delta_{T}^{-1}\big),\quad\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\eta_{t}\|\leq C\,\Delta_{T}\log\big(1+\Delta_{T}^{-1}\big)$
where $\Delta_{T}=\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\Phi(\mu_{t-1})\|$ with the convention that $\Phi(\mu_{0})=\eta_{1}$.
Proof.
See Section 5.4. ∎
Remark 5.
Let $(\mu_{t})_{t\geq 1}$ be a sequence in $\mathcal{P}(\mathbb{R})$ such that, for all $t\geq 1$, we have $\mu_{t+1}=\Phi_{t}(\mu_{t})$ for some mapping $\Phi_{t}:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$. Then, informally speaking, Lemma 1 implies that if $\Phi_{t}$ is close to $\Phi$ for all $t\in\{1,\dots,T\}$ then the estimation error $\|\mu_{t}-\eta_{t}\|$ will be small for all $t\in\{1,\dots,T\}$. This message conveyed by Lemma 1 is very intuitive, but, to the best of our knowledge, has not been proved to be true before.
4.2 Proof of Theorem 1
Proof.
In what follows we let $T\geq 1$ be fixed and we use the convention that $\hat{\eta}^{N}_{0}M=\eta_{1}$. Then, by Lemma 1, to obtain the result of Theorem 1 it suffices to control the behaviour of $\sup_{t\in\{1\,\dots,T\}}\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|$ as a function of $N$.
To this aim, remark first that, with $\mathbb{P}$-probability one and for all $t\geq 1$ and all $N\geq 1$, the function $F_{\hat{\eta}^{N}_{t-1}M}:\mathbb{R}\rightarrow(0,1)$ is continuous. Therefore, by the Dvoretzky-Kiefer-Wolfowitz inequality (see Massart,, 1990, Corollary 1), for all $(\gamma,\kappa)\in(0,1)$ we have
$\sup_{t\in\{1,\dots,T\}}\mathbb{P}\Big(\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|\leq\kappa\big|\mathcal{F}_{t-1}^{N}\Big)\geq 1-\gamma,\quad\forall N\geq N_{\kappa,\gamma}:=\frac{\log(2/\gamma)}{2\kappa^{2}}$ (15)
with the $\sigma$-algebra $\mathcal{F}_{t}^{N}$ as defined in the proof of Proposition 2 for all $t\geq 0$ and all $N\geq 1$. Using (15), we can conclude that
$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|\leq\kappa\Big)\geq(1-\gamma)^{T},\quad\forall N\geq N_{\kappa,\gamma},\quad\forall(\gamma,\kappa)\in(0,1).$
We now let $q\in(0,1)$. Then, by applying the latter result with $\gamma=1-(1-q)^{1/T}$, we obtain that
$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|\leq\kappa\Big)\geq 1-q,\quad\forall N\geq N_{\kappa,1-(1-q)^{1/T}}.$ (16)
By Bernoulli’s inequality (see e.g. Li and Yeh,, 2013), we have $1-(1-q)^{1/T}\geq T^{-1}q$ and thus
$N_{\kappa,1-(1-q)^{1/T}}\leq\frac{\log(2)+\log(T/q)}{2\kappa^{2}}\leq\frac{\log(2)+\log(T/q)}{\kappa^{2}}.$
Together with (16), this shows that
$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\hat{\eta}_{t-1}M\|\leq N^{-1/2}\sqrt{\log(2)+\log(T/q)}\Big)\geq 1-q$
and the result of Theorem 1 follows from Lemma 1. ∎
4.3 Proof of Theorem 2
4.3.1 Preliminary result: A discrepancy bound for sampling from univariate mixtures
The proof of the following lemma is inspired by that of Hlawka and Mück, (1972, “Statz 2”).
Lemma 2.
Let $x_{1},\dots,x_{M}$ be $M$ points in $\mathbb{R}$ for some $M\geq 1$, and for all $x\in\{x_{1},\dots,x_{M}\}$ let $\mu_{x}\in\mathcal{P}(\mathbb{R})$. In addition, let $\pi_{M}(\mathrm{d}x)=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\in\mathcal{P}(\mathbb{R})$ and $\mu=\sum_{m=1}^{M}(W_{m}\delta_{\{x_{m}\}}\otimes\mu_{x_{m}})\in\mathcal{P}(\mathbb{R}^{2})$. Lastly, let $\varphi_{\mu}:(0,1)^{2}\rightarrow\mathbb{R}^{2}$ be defined by
$\varphi_{\mu}(u)=\Big(z_{1},F^{-1}_{\mu_{x_{1}}}(u_{2})\Big),\quad z_{1}=F^{-1}_{\pi_{M}}(u_{1}),\quad u=(u_{1},u_{2})\in(0,1)^{2}$
and let $\{u^{n}\}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ for some $N\geq 1$. Assume that, for all $a\in\mathbb{R}$, the mapping $\{x_{1},\dots,x_{M}\}\ni x\mapsto F_{\mu_{x}}(a)$ is monotone and let $\mathcal{B}_{2}=\big\{\prod_{i=1}^{2}(-\infty,a_{i}],\,(a_{1},a_{2})\in\mathbb{R}^{2}\big\}$. Then,
Proof.
Letting $j_{B}=\max\{i\in\{1,\dots,M\}:x_{i}\leq a_{1}\}$, and using the convention that $W_{0}=0$, we note that
$\begin{split}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{G(B)}(u^{n})-\lambda_{2}(G(B))&=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})+\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\end{split}$ (20)
and we start by studying the first term on the r.h.s. of (20).
is such that $Q(\xi_{1})=(a,b]\subset(0,1]^{2}$ for some $a,b\in[0,1]^{2}$. Therefore, noting that $\cup_{\xi_{1}\in I}Q(\xi_{1})=U_{1}$, we obtain that
$\begin{split}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})\Big|\leq\sum_{\xi_{1}\in I}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{Q(\xi_{1})}(u^{n})-\lambda_{2}(Q(\xi_{1}))\Big|&\leq 4|I|D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)\\ &\leq 4L\,D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)\end{split}$ (21)
$\sup_{(a,b]\subset(0,1]^{2}}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{(a,b]}(u^{n})-\lambda_{2}\big((a,b]\big)\Big|\leq 4D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big).$ (22)
If $\mathsf{U}_{1}=\emptyset$ then $\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})=0$ and thus (21) also holds in this case.
We now study the second term on the r.h.s. of (20). To this end we note first that we can cover $U_{1}^{\prime}$ with sets in $\mathsf{U}_{2}$, and thus
By using this latter result, we obtain that
where the second inequality holds by (22).
$\sup_{B\in\mathcal{B}_{2}}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big(\varphi_{\mu}(u^{n})\big)-\mu(B)\Big|\leq 12LD^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)+2L^{-1}$ (24)
and the result of the lemma follows by applying (24) with $L=\big\lceil\big\{6D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)\big\}^{-1/2}\big\rceil$. ∎
4.3.2 Proof of the theorem
Proof of Theorem 2.
By Lemma 2, for all $N\geq 1$ and using the convention that $\hat{\eta}^{N}_{0}M=\eta_{1}$, we have $\sup_{t\geq 1}\|\eta_{t}^{N}-\hat{\eta}_{t-1}^{N}M\|\leq\sqrt{238}D^{*}_{2}\big(\{U_{t}^{n}\}_{n=1}^{N}\big)$ and the result of Theorem 2 then follows from Lemma 1 and (13). ∎
4.4 Proof of Corollary 1
Proof.
In what follows we let $(\kappa,q)\in(0,1)$ be arbitrary and $\delta_{N,T,q}=N^{-1}\big(1+\log(T/q)\big)$ for all integers $N\geq 1$ and $T\geq 1$. In addition, we let $C_{\kappa}\in(\exp(2),\infty)$ so that the inequality $\delta_{N,T,q}^{-1/2}\geq C_{\kappa}^{1/2}\kappa^{-1}\geq\exp(1)$ holds for all $N\geq C_{\kappa}\kappa^{-2}\big(1+\log(T/q)\big)$. Then, using the fact that function $f(x)=x^{-1}\log(1+x)$ is non-increasing on $x\in(\exp(1),\infty)$, it follows that
$N$ $\geq C_{\kappa}\kappa^{-2}\big(1+\log(T/q)\big)\implies\delta_{N,T,q}^{1/2}\log\big(1+\delta_{N,T,q}^{-1/2}\big)\leq C_{\kappa}^{-1/2}\kappa\log\big(1+C_{\kappa}^{1/2}\kappa^{-1}\big)$
where, with the constant $\bar{C}_{1}\in(0,1)$ as in the statement of Theorem 1,
$\bar{C}_{1}C_{\kappa}^{-1/2}\kappa\log\big(1+C_{\kappa}^{1/2}\kappa^{-1}\big)\leq\kappa\Leftrightarrow\frac{\log\big(1+C_{\kappa}^{1/2}\kappa^{-1}\big)}{C_{\kappa}^{1/2}}\leq\frac{1}{\bar{C}_{1}}.$ (25)
To proceed further we let $D\in(1,\infty)$ and $f_{D}:(0,\infty)\rightarrow\mathbb{R}$ be defined by
$f_{D}(x)=\frac{\log\big(1+xDv(x)\big)}{Dv(x)},\quad v(x):=\exp(1)+\log(1+x),\quad x\in(0,\infty).$
Simple calculations show that the function $x\mapsto\log(1+x)/\log(x)$ is non-increasing on the interval $(1,\infty)$. Therefore, noting that, since $D\geq 1$, we have $xDv(x)\geq\exp(1)>1$ for all $x\geq 1$, it follows that
$\begin{split}f_{D}(x)=\frac{\log\big(xDv(x)\big)}{Dv(x)}\frac{\log\big(1+xDv(x)\big)}{\log\big(xDv(x\big))}\leq\frac{\log\big(xDv(x)\big)}{Dv(x)}\log\big(1+\exp(1)\big),\quad\forall x\geq 1\end{split}$ (26)
where
$\frac{\log\big(xDv(x)\big)}{Dv(x)}\leq\frac{1}{D}\bigg(1+\frac{\log(D)}{\exp(1)}+\sup_{z\geq 1}\frac{\log\big(v(x))}{v(x)}\bigg),\quad\forall x\geq 1.$ (27)
Consequently, by (26)-(27), for any $\epsilon\in(0,\infty)$ there exists a constant $D_{\epsilon}\in(1,\infty)$ such that we have $\sup_{x\in[1,\infty]}f_{D}(x)\leq\epsilon$ for all $D\in[D_{\epsilon},\infty)$.
By using this latter result, and assuming without loss of generality that $\bar{C}_{1}>1$, it follows that for any constant $D\geq D_{1/\bar{C}_{1}}\geq 1$ the second inequality in (25) holds for $C_{\kappa}=D^{2}\big(\exp(1)+\log(1+\kappa^{-1})\big)^{2}>\exp(2)$. The result of the corollary then follows from Theorem 1. ∎
4.5 Proof of Corollary 2
Proof.
The result of Corollary 2 easily follows from Theorem 2 and the calculations done in the proof of Corollary 1. The details are therefore omitted to save space. ∎
5 Proof of Proposition 1 and proof of Lemma 1
5.1 Additional notation
@@
For a function $f\in\mathcal{C}^{1}(\mathbb{R})$ we let $V(f)=\int_{\mathbb{R}}|f^{\prime}(x)|\mathrm{d}x$ with the convention that $V(f)=0$ if $f^{\prime}(x)=0$ for all $x\in\mathbb{R}$. Abusing notation, for a function $f:\mathbb{R}\rightarrow\mathbb{R}$ continuously differentiable on some non-empty interval $(a,b)$ we let $V(\mathbbm{1}_{(a,b]}f)=V(\mathbbm{1}_{(a,b)}f)=\int_{a}^{b}|f^{\prime}(x)|\mathrm{d}x$. Next, for all $(z,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $M_{r,s^{2}}(z,\mathrm{d}x)=\mathcal{N}(rz,s^{2})$ and $G_{s}(z)=e^{-sz^{2}/2}$, and below $\varphi(\cdot)$ denote the probability density function of the $\mathcal{N}(0,1)$ distribution. Moreover, for all $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $h_{a,r,s}:\mathbb{R}\rightarrow[-1,1]$ be defined by
$h_{a,r,s}(x)=\Phi\bigg(\frac{a-rx}{s}\bigg)-\Phi\bigg(\frac{a}{s}\bigg),\quad x\in\mathbb{R}$ (28)
while, for any bounded function $H:\mathbb{R}\rightarrow(0,\infty)$, we let $\Psi_{H}:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be defined by
@@
5.2 Preliminary results
5.2.1 A simple but useful technical lemma